### v2

# AFD-SLU: Adaptive Feature Distillation for Spoken Language Understanding

Abstract
Spoken Language Understanding (SLU) is a core component of conversational systems, enabling machines to interpret user utterances. Despite its importance, developing effective SLU systems remains challenging due to the scarcity of labeled training data and the computational burden of deploying Large Language Models (LLMs) in real-world applications. To further alleviate these issues, we propose an Adaptive Feature Distillation framework that transfers rich semantic representations from a General Text Embeddings (GTE)-based teacher model to a lightweight student model. Our method introduces a dynamic adapter equipped with a Residual Projection Neural Network (RPNN) to align heterogeneous feature spaces, and a Dynamic Distillation Coefficient (DDC) that adaptively modulates the distillation strength based on real-time feedback from intent and slot prediction performance. Experiments on the Chinese profile-based ProSLU benchmark demonstrate that AFD-SLU achieves state-of-the-art results, with 95.67% intent accuracy, 92.02% slot F1 score, and 85.50% overall accuracy.
Index Terms—  Knowledge Distillation, Spoken Language Understanding, General Text Embeddings, Feature Knowledge
Fig. 1: An overview of our model AFD-SLU.
1 Introduction
Spoken Language Understanding (SLU) aims to transform human utterances into structured semantic representations and typically involves two core and closely related tasks: intent detection and slot filling [15]. As a fundamental component of Natural Language Processing (NLP), SLU supports a broad range of applications, such as voice assistants, customer service systems, and smart devices.
Early research in SLU primarily focused on capturing the inherent correlation between intent detection and slot filling [2, 3, 14, 10, 8]. To further enhance model performance, subsequent work has explored knowledge distillation techniques leveraging pre-trained language models [13]. Li et al. [5] proposed a multi-teacher framework that transfers knowledge from multiple pre-trained models under limited supervision. However, task-level distillation are tailored to specific domains or languages, exhibiting limited cross-domain adaptability and generalization. With the rapid advancement of Large Language Models (LLMs), recent studies have begun to harness their strong generalization and few-shot capabilities for SLU [19, 11, 16]. Notably, Yin et al. [18] reformulated slot filling as an entity recognition task and proposes a ”Chain of Intent” mechanism to support multi-intent understanding within an LLM-based framework. Despite their promising performance, LLMs entail significant memory and computational costs, which poses challenges for their deployment in real-world, resource-constrained SLU applications [1, 4].
These challenges are also prevalent in the Chinese. To address data scarcity, Xu et al. [17] introduced ProSLU, a high-quality profile-based benchmark designed to address data scarcity. Building upon this resource, recent studies have explored the integration of user profile information to improve SLU performance. Thinh et al. [7] proposed a joint model with slot-to-intent attention, while Teng et al. [12] developed a heterogeneous graph attention network that leverages profile features. Although these models perform well on ProSLU benchmark, they still encounter performance bottlenecks due to the dataset’s limited scale.
To address these limitations, we propose Adaptive Feature Distillation for Spoken Language Understanding (AFD-SLU), a lightweight framework for dynamic feature knowledge distillation. It comprises three components: a teacher model, a student model, and a dynamic adapter. The teacher model is a General Text Embeddings (GTE) model [6] derived from a high-performance LLM, which provides generalizable semantic features while avoiding the heavy computational cost of direct LLM inference. To the best of our knowledge, this is the first work to employ GTE as a teacher model for SLU. The student model adopts a joint architecture for intent detection and slot filling, while the dynamic adapter integrates a Residual Projection Neural Network (RPNN) to align feature spaces and a Dynamic Distillation Coefficient (DDC) to adaptively adjust the distillation weight. Experiments on the ProSLU benchmark demonstrate that AFD-SLU achieves state-of-the-art performance. Ablation studies verify the effectiveness of each module, and further experiments show the adaptability of our framework to different teacher models. All codes for this work are publicly available at https://github.com/XieyRock/AFD-SLU.
2 APPROACH
This section presents the architecture of our AFD-SLU, as shown in Figure 1. The framework consists of three main components: a teacher model that provides token-level semantic embeddings (§2.1), a student model for joint intent detection and slot filling (§2.2), and a dynamic adapter for feature alignment and adaptive training (§2.3).
2.1 Student Model
The student model adopts a joint SLU architecture designed for Chinese linguistic characteristics and optimized for efficient inference. In this work, we adopt a Bidirectional Long Short-Term Memory (BiLSTM)-based joint SLU model [7, 12] as the student model, which is currently one of the most commonly used architectures for joint models. Given an input sequence $X=\{{{x}_{1}},{{x}_{2}},\cdots,{{x}_{n}}\}$ with n tokens, a BiLSTM followed by a self-attention layer is applied to capture sequential and contextual information. An attention pooling layer is then applied to aggregate these features, yielding a representation $e_{i,j}^{s}\in\mathbb{R}^{d_{b}\times d_{es}}$, where $d_{b}$ is the batch size and $d_{es}$ denotes the dimensionality of the student model’s text embeddings. Here, $i$ refers to the $i$-th sample in the batch and $j$ denotes the $j$-th token in the sequence.
2.2 Teacher Model
We adopt the GTE model [6] as the teacher model. During the entire training process, the teacher model is solely responsible for generating sentence embeddings and providing supervision signals to the student model, with its parameters kept frozen. The generation of sentence embeddings involves the following three steps:
Utterance Encoding: Given a user utterance $X=\{x_{1},x_{2},\cdots,x_{n}\}$, identical to the one used in the student model, it is first tokenized into a sequence $X\in\mathbb{Z}^{L}$ using a pre-trained tokenizer. To conform to the input constraints of the student model, the token sequence is padded to a fixed length $L$, resulting in a standardized input tensor. Meanwhile, an attention mask $m\in\{0,1\}^{L}$ is constructed using a standard masking algorithm to indicate valid token positions.
Feature Extraction: We extract the outputs from the last four hidden layers of the teacher model, denoted as $H=[h_{1},h_{2},h_{3},h_{4}]\in\mathbb{R}^{4\times L\times d_{et}}$, where $d_{et}$ is the dimensionality of the teacher model’s text embeddings. Each $h_{k}$ corresponds to the hidden representation from the $k$-th last layer.
Embedding Generation: The four hidden states are averaged to obtain a unified feature representation. This representation is then passed through mean pooling over the valid tokens (as indicated by the attention mask) to derive the final sentence embeddings. The computation is defined as:
$e_{i,j}^{T}=\frac{\sum\limits_{l=1}^{L}\bar{h}_{l}\cdot m_{l}}{\sum\limits_{l=1}^{L}m_{l}}$ (1)
Here, $\bar{h}_{l}$ represents the averaged hidden state at position $l$ across the last four layers of the teacher model. The representation $e_{i,j}^{T}\in\mathbb{R}^{d_{b}\times d_{et}}$ denotes the teacher model’s embeddings of the $j$-th token in the $i$-th sample.
2.3 Dynamic Adapter
2.3.1 Residual Projection Neural Network
In our model, the Residual Projection Neural Network (RPNN) plays a critical role in aligning feature representations between the teacher and student model. Its primary goal is to project the lower-dimensional embedding space of the student model into the higher-dimensional space of the teacher model through a nonlinear transformation, while preserving essential information from the original features.
Let the text embeddings output from the student model be denoted as $e_{i,j}^{s}\in\mathbb{R}^{d_{b}\times d_{es}}$, and the corresponding embeddings from the teacher model as $e_{i,j}^{T}\in\mathbb{R}^{d_{b}\times d_{et}}$, where $d_{es}\neq d_{et}$. The RPNN first projects $e_{i,j}^{s}$ into a higher-dimensional hidden space using a linear layer, followed by a GELU activation and Layer Normalization (LN):
$h_{1}=\text{LN}_{1}\left(\text{GELU}(W_{1}e_{i,j}^{s}+b_{1})\right)$ (2)
where $W_{1}\in\mathbb{R}^{d_{es}\times d_{h}}$, $b_{1}\in\mathbb{R}^{d_{h}}$, and $d_{h}=4d_{es}$.
To further enhance representational capacity, we introduce a residual transformation module, denoted as ResidualBlock, which is shown on the right of Figure 1. This module refines $h_{1}$ through a two-layer feed-forward network with a skip connection, followed by another layer normalization:
$h_{2}=\text{LN}_{2}\left(h_{1}+W_{3}\cdot\text{GELU}(W_{2}h_{1}+b_{2})+b_{3}\right)$ (3)
where $W_{2}\in\mathbb{R}^{d_{h}\times d_{h}}$, $W_{3}\in\mathbb{R}^{d_{h}\times d_{es}}$, $b_{2}\in\mathbb{R}^{d_{h}}$, and $b_{3}\in\mathbb{R}^{d_{es}}$. The residual connection facilitates gradient flow and preserves information from earlier layers.
Finally, the refined features are projected into the teacher model’s embeddings space:
$e_{i,j}^{S}=W_{4}h_{2}+b_{4}$ (4)
where $W_{4}\in\mathbb{R}^{d_{es}\times d_{et}}$ and $b_{4}\in\mathbb{R}^{d_{et}}$.
2.3.2 Dynamic Distillation Coefficient
Fixed-weight strategies often fail to balance the distillation loss and the task loss effectively. They cannot reflect the changing learning capacity of the student model during training. This limits the efficiency of knowledge transfer.
We introduce a Dynamic Distillation Coefficient (DDC) that adaptively balances distillation and task losses during training. The distillation loss is computed as the Mean Squared Error (MSE) between the token-level embeddings of the teacher and student models:
$\mathcal{L}_{\text{distill}}=\frac{1}{d_{b}}\sum_{i=1}^{d_{b}}\sum_{j=1}^{d_{et}}\left(e_{i,j}^{T}-e_{i,j}^{S}\right)^{2}$ (5)
The dynamic coefficient $\lambda$ modulates the relative contribution of the distillation and task losses throughout training:
$\lambda=\lambda_{\text{final}}+(\lambda_{\text{initial}}-\lambda_{\text{final}})\left(1+\cos\left(\frac{e}{E}\pi\right)\right)$ (6)
where $\lambda_{\text{initial}}$ and $\lambda_{\text{final}}$ denote the initial and final weights, $e$ is the current epoch, and $E$ is the total number of epochs. This cosine annealing schedule gradually reduces the influence of the teacher model, enabling the student model to increasingly focus on task-specific learning in the later stages of training.
$\mathcal{L}_{\text{task}}=\alpha\mathcal{L}_{\text{ID}}+(1-\alpha)\mathcal{L}_{\text{SF}}$ (7)
where $\mathcal{L}_{\text{ID}}$ and $\mathcal{L}_{\text{SF}}$ denote the intent loss and the slot loss.
The overall training objective is formulated as:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{task}}+\lambda\mathcal{L}_{\text{distill}}$ (8)
3 EXPERIMENTS
3.1 Experimental Setup
Dataset and Evaluation metrics. To evaluate the effectiveness of the proposed framework, we conduct experiments on the ProSLU dataset, a Chinese spoken language understanding benchmark that incorporates user profile information. The dataset contains 4,196 training utterances, 522 validation utterances, and 531 test utterances. Each utterance is accompanied by four user profile vectors and four context-aware vectors, and the annotations cover 14 intents and 99 slots. Model performance is assessed using three standard metrics: slot F1 score, intent accuracy, and overall accuracy.
Implementation Details. We train our model on a NVIDIA RTX 4090D GPU equipped with 48 GB of memory. We optimize the model using the Adam and train it for 50 epochs, applying a dropout rate of 0.4. The dynamic parameter $\lambda$ was selected based on comparative validation. Its initial and final values were set to $\lambda_{\text{initial}}=0.1$ and $\lambda_{\text{final}}=0.7$, respectively. We evaluate our method with both JPIS [7] and PRO-HAN [12] serving individually as student models, and employ the gte_Qwen2-1.5B-instruct [6] as teacher model. For JPIS, we use a batch size of 1 and a learning rate of $2\times 10^{-4}$; for PRO-HAN, the batch size is 64 and the learning rate is $1\times 10^{-3}$.
3.2 Main Results
We apply our AFD-SLU framework to the above student–teacher settings. As shown in Table 1, integrating AFD improves performance across all evaluation metrics. For JPIS, applying AFD yields absolute gains of 2.25% in intent accuracy, 2.75% in slot F1 score, and 3.20% in overall accuracy. Similarly, PRO-HAN-AFD achieves 95.67% intent accuracy and 92.02% slot F1 score, outperforming the original model by 3.58% and 2.01%, respectively, while maintaining comparable overall accuracy. These results demonstrate that AFD-SLU significantly enhances model performance under low-resource conditions, while keeping the model parameters and the training data scale nearly unchanged. The consistent gains confirm that our method effectively transfers semantic knowledge from GTE models to student models, improving both intent detection and slot filling without increasing inference overhead.
Table 1: Main results on the ProSLU dataset. Models with “+AFD” denote variants enhanced with adaptive feature distillation. indicates our reproduced results.
Model Intent (Acc) Slot (F1) Overall (Acc) Params SF-ID [2] 83.24 73.70 68.36 <<1M Slot-Gated [3] 83.24 74.18 69.11 <<1M Bi-Model [14] 81.92 77.76 73.45 $\sim$10M AGIF [10] 81.54 80.57 74.95 $\sim$1M Stack-Propagation [8] 83.99 81.08 78.91 $\sim$10M General-SLU [17] 85.31 83.27 79.10 $\sim$1M GL-GIN [9] 85.69 82.70 79.28 $\sim$1M JPIS [7] 87.95 85.76 82.30 $\sim$2M JPIS-AFD (Ours) 90.20 88.51 85.50 $\sim$2M PRO-HAN [12] 92.09 90.01 85.31 $\sim$40M PRO-HAN-AFD (Ours) 95.67 92.02 85.50 $\sim$40M
Table 2: Ablation study results using JPIS as the student model and gte_Qwen2-1.5B-instruct as the teacher model.
Model Intent (Acc) Slot (F1) Overall (Acc) JPIS-AFD 90.20 88.51 85.50 RPNN(Linear) 87.95 86.13 83.43 RPNN(Deep) 88.70 87.13 83.99 w/o DDC 87.59 85.96 83.24
3.3 Ablation study
We perform an ablation study to assess the contributions of key components in our framework. As shown in Table 2, RPNN (Linear) applies a single linear projection and results in performance drops of 2.01% in intent accuracy, 2.38% in slot F1 score, and 2.07% in overall accuracy compared to the JPIS-AFD model. Meanwhile, RPNN (Deep) introduces a deeper nonlinear projection network, also leads to reduced performance.This observation suggests that effective feature alignment benefits from a carefully balanced RPNN design, whereas both overly simple linear transformations and excessively complex projection networks are suboptimal. Similarly, excluding the DDC leads to reductions of 2.61%, 2.55%, and 2.26% in the same metrics. This highlights the importance of adaptively adjusting the distillation weight to balance the student’s intrinsic learning and the knowledge transferred from the teacher during training. These results confirm that both RPNN and DDC are essential to the effectiveness of our method, working together to significantly enhance performance on joint SLU tasks.
3.4 Analysis
We evaluated the impact of different GTE models used as teacher models, with results shown in Table 3. We can conclude that: Smaller or moderately sized teacher models that are well-aligned with the task, such as GTE-Qwen2-1.5B-instruct and Qwen3-Embedding-0.6B, are more effective in guiding the student model through knowledge distillation. These models achieve better performance across intent recognition, slot filling, and overall accuracy. In contrast, overly large models tend to overfit when trained on small datasets such as ProSLU, exhibiting high initial accuracy but subsequently suffering from overfitting, which leads to suboptimal generalization and degraded performance. Additionally, pre-trained models such as GTE-large-zh may lack sufficient task-specific understanding, which limits their effectiveness as teacher models. These findings emphasize the importance of selecting a teacher model that is not only powerful but also well-suited to the task and dataset scale.
Table 3: Impact of different general text embedding models used as teacher models on knowledge distillation performance, with JPIS as the student model.
Teacher Model Intent (Acc) Slot (F1) Overall (Acc) gte_large-zh 88.14 86.11 83.24 gte_Qwen2-1.5B-instruct 90.20 88.51 85.50 gte_Qwen2-7B-instruct 89.64 88.00 84.75 Qwen3-Embedding-0.6B 90.21 88.47 85.12 Qwen3-Embedding-4B 88.14 86.80 84.18 Qwen3-Embedding-8B 87.01 85.27 82.30
4 CONCLUSION
In this paper, we propose AFD-SLU, an Adaptive Feature Distillation framework aimed at enhancing Chinese SLU in low-resource scenarios. AFD-SLU leverages a powerful GTE model as a teacher to transfer extensive semantic knowledge to a lightweight student model. Extensive experiments on the ProSLU benchmark demonstrate that AFD-SLU achieves state-of-the-art performance, significantly outperforming baselines in intent detection, slot filling, and overall accuracy. Ablation studies validate the contribution of each component, and comparisons across teacher models reveal that models with larger parameter sizes do not necessarily lead to better student performance; selecting a teacher suited to the task is crucial. Future work may explore more effective data augmentation and model compression to further enhance practicality in real-world low-resource scenarios.